A parcel of lot is to be sold and tobe divided among the four heirs of the spouses carl smith and helen smith. the lenght of a rectangular lot is three times its width, if the lenght in increased by five meters and the width is decreased by two meters the original area of the lot will decrease by twenty one sq. meters. Find the dimension of the lot. Using the new area, what is the price of the lot if it was sold at $5000 per sq. meters? how much does each heir will receive if heirs one, two, three and four will be receiving 10%, 20%, 30% and 40% respectively?
Answer by Boreal(15235)

L=3W
Area is L*W=3W^2
New dimensions are (L+5)(W-2)=3W^2-21
But L still equals 3W
(3W+5)(W-2)+3W^2-21
3W^2-W-10=3W^2-21
-W=-11
W=11
L=33
original area is 363 m^2
new L is 38
new W is 9
new area is 342 m^2, which is 21 m^2 fewer.
at $5000/m^2, the lot gets $1,710,000.
Heir 1 gets $171,000
2---$342,000
3---$513,000
4---$684,000